# Functional Protein Design with Local Domain Alignment

### v2

@@ Abstract
The core challenge of de novo protein design lies in creating proteins with specific functions or properties, guided by certain conditions. Current models explore to generate protein using structural and evolutionary guidance, which only provide indirect conditions concerning functions and properties. However, textual annotations of proteins, especially the annotations for protein domains, which directly describe the protein’s high-level functionalities, properties, and their correlation with target amino acid sequences, remain unexplored in the context of protein design tasks. In this paper, we propose Protein-Annotation Alignment Generation (PAAG), a multi-modality protein design framework that integrates the textual annotations extracted from protein database for controllable generation in sequence space. Specifically, within a multi-level alignment module, PAAG can explicitly generate proteins containing specific domains conditioned on the corresponding domain annotations, and can even design novel proteins with flexible combinations of different kinds of annotations. Our experimental results underscore the superiority of the aligned protein representations from PAAG over 7 prediction tasks. Furthermore, PAAG demonstrates a nearly sixfold increase in generation success rate (24.7% vs 4.7% in zinc finger, and 54.3% vs 8.7% in the immunoglobulin domain) in comparison to the existing model. We anticipate that PAAG will broaden the horizons of protein design by leveraging the knowledge from between textual annotation and proteins.
1 Introduction
@@
Protein design Marshall et al. (2019) is a crucial task for its immense potential on drug discovery Sliwoski et al. (2014), enzyme engineering Sterner et al. (2008), immunongineering Swartz et al. (2012) and so on. The generation of proteins with specific properties, behaviors, or functions, such as optimizing the binding affinity to given molecules Houk et al. (2003) or incorporating a particular ion-binding site Regan (1995), is known as de novo protein design. This process presents a significant challenge due to the vast space of protein sequences and the complexity of protein functions. Recently, machine learning models have shown profound potential for protein design. The existing studies mostly rely on the structural Watson et al. (2023) or evolutionary information Alamdari et al. (2023a) as the guidance to design proteins. However, in many cases, these conditions can only offer indirect guidance towards the desired protein design targets to their inherent ambiguity. For example, the same protein sequence segment can be either act as receptors to regulate synaptic function Stathakis et al. (1997) or helpers to locate target proteins to specific subcellular locations Mohandas et al. (1995).
@@
Figure 1: (a) The example of property annotations (in bold) and domain annotations (in colors). (b). The illustration of annotation-guided protein design with PAAG. Given the input of textual description within immunoglobulin domain annotation, PAAG can generate the proteins containing immunoglobulin domain.
@@
For instance, as one of crucial functional domains of DNA/RNA binding proteins Cassandri et al. (2017), the zinc-finger domain naturally has many variants, such as C2H2 type, CCHC type and Zinc ribbon type. These variants exhibit significant differences in both structural and evolution features among them which is hard covered by structural and evolutionary conditions. On the contrary, the “zinc-finger” annotation from the protein database inherently provides a more effective means of describing the high-level knowledge span across both sequences and structures. Hence, we aim to investigate the following question: Is it possible to leverage such textual annotations to guide the delicate controllable protein design?
@@
Recently, several primary attempts have been made to leverage such textual annotations to guide the protein generation. Examples include extracting tags from labels to create unique tokens Madani et al. (2023), training an individual protein caption model to guide the diffusion generation process Ingraham et al. (2023), and incorporating the an overall text description through a global language-to-protein alignment model Liu et al. (2023b). However, current models cannot flexibly combine the different conditions and lack of the capability for fine-grained control, such as specifying the generation of particular domains.
@@
To fill these gaps, in this paper, we introduce a novel framework, Protein-Annotation Alignment Generation(PAAG), that enables annotation-guided protein design by aligning protein sequences with their textual annotations. Specifically, we first consider both property and domain annotations in proteins and design a multi-level alignment module to align the representations of sequences and annotations extracted from the existing encoders in both global and local level. For the generation task, PAAG utilizes an autoregressive decoder to generate protein sequences guided by the aligned representation of textual annotations. Additionally, PAAG employs an end-to-end training pipeline that joint training of alignment and generation tasks without freezing the parameters in sequence and text encoders. This joint training enhances the understanding of the complex and flexible annotation condition, resulting in improved guided generation. Figure 1 demonstrates an example of annotation-guided generation.
In experiments, we first investigate the quality of protein representations from PAAG by seven predictive downstream tasks. PAAG surpasses state-of-the-art benchmarks with an average relative improvement of 1.5%. Subsequently, three protein generation tasks are conducted to assess the capabilities of PAAG. In the case of unconditional generation, PAAG produces sequences exhibiting the highest degree of novelty while maintaining the distribution of natural proteins. For the two conditional protein design tasks, PAAG demonstrates a nearly threefold increase in generation success rate (24.7% vs 4.7% in zinc finger, and 54.3% vs 8.7% in the immunoglobulin domain) [^1] This is the success rate with quality threshold $e=1$, i.e., $\text{SR}_{1}$. in comparison to the existing model.
Our contributions are summarized as follows:
@@
  • To the best of our knowledge, we are the first to propose a new annotation-guided paradigm for protein design, which can simultaneously integrate local and global annotation information. Our proposed framework PAAG is the first approach that can generate proteins containing specific domains, guided by their corresponding annotations with high success rate.
@@
  • Comprehensive experiments on 7 predictive and 3 generative tasks showcase PAAG’s superiority compared to the exisiting methods. Notably, PAAG is not only capable of generating proteins that include a single domain, but it also successfully generates proteins adhering to the flexible conditions of multiple annotation combinations.
2 Preliminaries
2.1 Protein and Its Textual Annotations
The primary structure of a protein can be represented as an amino acid sequence $S=(s_{1},s_{2},\cdots,s_{n})$, where $s_{i}$ is the $i$-th amino acids chosen from 20 different amino acids which are represented as 20 characters.
@@ 2.1 Protein and Its Textual Annotations
Given a protein $S$, the annotation-sequence pair set $\mathcal{D}_{S}=\{(T_{k},S^{k}_{i:j})\}_{k=1}^{K_{S}}$ is constructed by extracting the correspondence between textual annotations and protein domains & properties from protein database, such as UniProtKB Consortium (2022), where $T_{k}$ is the textual annotations of the $k$-th domain $S^{k}_{i:j}$ of the protein and $K_{S}$ is the number of annotation-domain pairs of the protein $S$. For the property annotation, the corresponding domain is the entire sequence, i.e, $S_{1:|S|}$. This annotation-domain pair set $\mathcal{D}_{S}$ provides comprehensive knowledge about the protein $S$ and, therefore, plays an important role in training the protein generation model.
2.2 Encoders for Proteins and Annotations
We adopt a protein encoder (PE) to generate both the protein and its domain representations based on the amino acid sequence as $\bm{z}_{S}=f_{\text{PE}}(S)$, where $f_{\text{PE}}$ is the protein encoder and $\bm{z}_{S}$ is the embedding of $S$.
@@ 2.2 Encoders for Proteins and Annotations
For textual annotations, we generate their representations using the pre-trained language model $f_{\text{LM}}$ as $\bm{z}_{\mathcal{A}_{S}}=f_{\text{LM}}(G(\mathcal{A}_{S}))$, where $\mathcal{A}_{S}\subseteq\{T_{k}|(T_{k},S_{i:j}^{k})\in\mathcal{D}_{S}\}$ is a subset of the annotations in the annotation-domain pair set for a protein $S$, $G()$ is a template function which converts a set of annotations to a textual description, and $\bm{z}_{\mathcal{A}}$ is the embedding of the annotation set. Note that the annotation subset can cover all annotations in $\mathcal{D}_{S}$ or just one. If $\mathcal{A}_{S}$ covers only one annotation $T_{0}$, we can skip the template function and directly generate the embedding of $T_{0}$ using $f_{\text{LM}}$.
@@
We employ transformer-based protein encoder and text encoder as our base model and initialize them with pre-trained models. Specifically, we employ SciBERTBeltagy et al. (2019), which is pre-trained on computer science and biological datasets, to initialize the text encoder $f_{\text{LM}}$. The pre-trained protein encoder, ProtBERTElnaggar et al. (2021), is used to initialize the protein encoder $f_{\text{PE}}$.
@@
Figure 2: The overall framework of PAAG. The same parameters share the same color. PAAG contains three modules. (1) Protein & Annotation Encoding module encode the input protein sequence & domains and corresponding annotations to the embeddings. (2) Multi-level alignment module projects the protein and annotation embeddings into and employs Annotation-Protein Contrasive (APC) loss, Annotation-Domain Contrasive (ADC) loss and Annotation-Protein Matching (APM) loss to align them in a same latent space. (3) Conditional Protein Decoding accepts the annotation embedding as input and generate the protein sequence.
3 Methodology
@@
In this section, we propose a multi-modal framework, Protein-Annotation Alignment Generation(PAAG), which enables the flexible annotations-guided protein design. Figure 2 illustrates the overall framework of PAAG. In the following, we present the details of each component of PAAG.
3.1 Conditional Protein Decoding
@@
Our ultimate goal is generating functional proteins conditioned on a set of annotations $\mathcal{A}=\{T_{k}\}_{k=1}^{K}$. We will first employ template function $G$ to translate these annotations as textual descriptions and then utilize a language model to extract the representation of the annotation set $\bm{z}_{\mathcal{A}}=f_{\text{LM}}(G(\mathcal{A}))$. Then the protein sequence is generated using a decoder $f_{D}$ based on this representation as $S=f_{\text{D}}(\bm{z}_{\mathcal{A}})$.
Protein Decoder. We adopt an auto-regressive protein decoder $f_{\text{D}}$ that can receive the condition from the language model. Specifically, the decoder will first process the protein sequence through causal attention layers to enable auto-regressive generation, then integrate the information from annotations via cross-attention layers followed by feed-forward layers. Note that the causal attention layers are initialized using the same weights as the self-attention layers in the protein encoder and the feed-forward layers share the same parameters as the protein encoder. Here, the parameter-sharing mechanism enables higher training efficiency Li et al. (2023). The cross-attention layers are randomly initialized and trained from scratch.
$\mathcal{L}_{PM}=-\sum_{i=1}^{l}\log{p(S_{i}|S_{<i};\mathcal{A})},$ (1)
3.2 Multi-level Protein and Annotation Alignment
@@
$s(\mathcal{A},S)=\frac{\langle\text{Proj}_{\mathcal{A}}(z_{\mathcal{A}}),\text{Proj}_{S}(\bm{z}_{S})\rangle}{\|\text{Proj}_{\mathcal{A}}(\bm{z}_{\mathcal{A}})\|\|\text{Proj}_{S}(\bm{z}_{S})\|},$ (2)
@@ 3.2.1 Local Alignment
Annotation-Domain Contrastive(ADC) Loss: We adopt InfoNCE loss as the ADC loss to align the representations of local annotation $\bm{z}_{T}$ and functional domain $\bm{z}_{S_{i:j}}$ as
$\begin{split}\mathcal{L}_{\mathrm{ADC}}=-\frac{1}{K_{S}}\sum_{k=1}^{K_{S}}&\log\frac{\exp(s({S^{k}_{i:j}},{T_{k}})/\tau)}{\sum_{n=1}^{N}\exp(s({S^{n-}_{i:j}},{T_{n}})/\tau)}\end{split},$ (3)
@@
3.2.2 Global Alignment
To enable global alignment, we utilize template function $G(\cdot)$ to construct protein-level textual description and form the positive protein-description pairs as $\{(G(\mathcal{A}_{S}),S)\}$. Since multiple annotations and the entire protein will be more complex, to enlarge the number of negative samples for global alignment, we follow the setting of MOCO He et al. (2020) to construct momentum encoders $f_{m}$ as
$f_{m}\leftarrow mf_{m}+(1-m)f.$ (4)
where the encoder $f$ can be either protein encoder $f_{\text{PE}}$ or text encoder $f_{\text{LM}}$, and $m$ is the momentum hyperparameter. We follow implementation details in Li et al. (2021) and Li et al. (2022) to construct the momentum encoders for both encoders.
The momentum encoders extract consistent features to increase the number of negative samples, and the dynamic dictionaries will store these features.
@@ 3.2.2 Global Alignment
$p_{m}^{\mathrm{s2a}}(S)=\frac{\exp(s(S,A_{m})/\tau)}{\sum_{m=1}^{M}\exp(s(S,A_{m})/\tau)},$ (5)
$p_{m}^{\mathrm{a2s}}(A)=\frac{\exp(s(A,S_{m})/\tau)}{\sum_{m=1}^{M}\exp(s(A,S_{m})/\tau)},$ (6)
@@
$\vspace{-1ex}\begin{split}\mathcal{L}_{\mathrm{APC}}=\frac{1}{2}\mathbb{E}_{(S,A)\sim D}\big{[}\mathrm{H}(\vec{y}^{\mathrm{~{}s2a}}(S),\vec{p}^{\mathrm{~{}s2a}}(S))\\ +\mathrm{H}(\vec{y}^{\mathrm{~{}a2s}}(A),\vec{p}^{\mathrm{~{}a2s}}(A))\big{]}\end{split}.$ (7)
Annotation-Protein Matching (APM) Loss. Inspired by  Li et al. (2022), we introduce a multimodal encoder $f_{\text{ME}}$, which integrates the representation of annotations and protein as $\bm{z}_{S,\mathcal{A}}^{M}=f_{\text{ME}}(S,G(\mathcal{A}))$, to identify whether the given pairs of protein description $G(\mathcal{A})$ and protein $S$ are matched or not. Specially, the multimodal encoder $f_{\text{ME}}$ is transformer-based and shares the parameters of self-attention and feed-forward layers with $f_{\text{PE}}$ and has additional cross-attention layers between self-attention layers and feed-forward layers to integrate the text information. The cross-attention layers share the same cross-attention parameters as the decoder.
@@
$\vspace{-1ex}\mathcal{L}_{\mathrm{APM}}=\mathbb{E}_{(A,S)}\mathrm{H}(\vec{y}^{\textrm{APM}}(A,S),\vec{p}^{\textrm{APM}}(A,S)).$ (8)
3.3 Training Objectives
In the training of PAAG, we optimize four objective functions. These functions are designed to align the representations between annotations and protein sequences, integrate the two modalities, and reconstruct the protein sequence. In contrast to ProteinDT Liu et al. (2023a), which splits the training process into three separate stages, PAAG jointly optimizes these four objective functions in an end-to-end manner.
The overall pretraining objective of PAAG is :
$\vspace{-1ex}\mathcal{L}=\mathcal{L}_{\mathrm{ADC}}+\mathcal{L}_{\mathrm{APC}}+\mathcal{L}_{\mathrm{APM}}+\mathcal{L}_{\mathrm{PM}}.$ (9)
3.4 Annotation-guided Protein Design

### v3

# Annotation-guided Protein Design with Multi-Level Domain Alignment

@@ Abstract.
The core challenge of de novo protein design lies in creating proteins with specific functions or properties, guided by certain conditions. Current models explore to generate protein using structural and evolutionary guidance, which only provide indirect conditions concerning functions and properties. However, textual annotations of proteins, especially the annotations for protein domains, which directly describe the protein’s high-level functionalities, properties, and their correlation with target amino acid sequences, remain unexplored in the context of protein design tasks. In this paper, we propose Protein-Annotation Alignment Generation (PAAG), a multi-modality protein design framework that integrates the textual annotations extracted from protein database for controllable generation in sequence space. Specifically, within a multi-level alignment module, PAAG can explicitly generate proteins containing specific domains conditioned on the corresponding domain annotations, and can even design novel proteins with flexible combinations of different kinds of annotations. Our experimental results underscore the superiority of the aligned protein representations from PAAG over 7 prediction tasks. Furthermore, PAAG demonstrates a significant increase in generation success rate (24.7% vs 4.7% in zinc finger, and 54.3% vs 22.0% in the immunoglobulin domain) in comparison to the existing model. We anticipate that PAAG will broaden the horizons of protein design by leveraging the knowledge from between textual annotation and proteins.
Annotation-guided protein design, multi-modality alignment
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Computational biology
1. Introduction
@@
Protein design (Marshall et al., 2019) is a crucial task for its immense potential on drug discovery (Sliwoski et al., 2014; Wu et al., 2024), enzyme engineering (Sterner et al., 2008), immunongineering (Swartz et al., 2012) and so on. The generation of proteins with specific properties, behaviors, or functions, such as optimizing the binding affinity to given molecules (Houk et al., 2003; Zhao et al., 2023) or incorporating a particular ion-binding site (Regan, 1995), is known as de novo protein design. This process presents a significant challenge due to the vast space of protein sequences and the complexity of protein functions. Recently, machine learning models have shown profound potential for protein design. The existing studies mostly rely on the structural (Watson et al., 2023) or evolutionary information (Alamdari et al., 2023a) as the guidance to design proteins. However, in many cases, these conditions can only offer indirect guidance towards the desired protein design targets to their inherent ambiguity. For example, the same protein sequence segment can be either act as receptors to regulate synaptic function (Stathakis et al., 1997) or helpers to locate target proteins to specific subcellular locations (Mohandas et al., 1995).
@@
Figure 1. (a) The example of property annotations (in bold) and domain annotations (in colors). (b). The illustration of annotation-guided protein design with PAAG. Given the input of textual description within immunoglobulin domain annotation, PAAG can generate the proteins containing immunoglobulin domain.
In addition to the structural and evolutionary information, the current protein dataset, such as Swiss-Prot (Bairoch and Apweiler, 2000) and UniProtKB (Consortium, 2022), contains rich textual annotations derived from wet laboratory experiments and literature. Figure 1 illustrates an example of the textual annotations on the zinc-finger protein. Generally, these annotations can be categorized into property annotation and domain annotation. Property annotation represents a piece of text that depicts the global property of proteins, such as protein names, number of amino acids, subcellular localization (Almagro Armenteros et al., 2017) and thermostability (Dallago et al., 2021). Conversely, domain annotation pertains to the knowledge derived from the local domain (Drenth et al., 1968) of proteins, which is a subregion of amino acid sequence that is self-stabilizing and represents certain structural and functional aspects of the protein. These annotations provide both coarse and fine-grained information regarding protein’s functions, properties, and interactions, thereby encompassing knowledge with the potential to guide the generation and design of novel proteins.
@@
For instance, as one of crucial functional domains of DNA/RNA binding proteins (Cassandri et al., 2017), the zinc-finger domain naturally has many variants, such as C2H2 type, CCHC type and Zinc ribbon type. These variants exhibit significant differences in both structural and evolutionary features among them which is hard covered by structural and evolutionary conditions. On the contrary, the “zinc-finger” annotation from the protein database inherently provides a more effective means of describing the high-level knowledge span across both sequences and structures. Hence, we aim to investigate the following question: Is it possible to leverage such textual annotations to guide the delicate controllable protein design?
@@
Recently, several primary attempts have been made to leverage such textual annotations to guide the protein generation. Examples include training an individual protein caption model to guide the diffusion generation process (Ingraham et al., 2023), and incorporating the an overall text description through a global language-to-protein alignment model (Liu et al., 2023b). However, current models cannot flexibly combine the different conditions and lack of the capability for fine-grained control, such as specifying the generation of particular domains.
@@
To fill these gaps, in this paper, we introduce a novel framework, Protein-Annotation Alignment Generation(PAAG), that enables annotation-guided protein design by aligning protein sequences with their textual annotations. Specifically, we first consider both property and domain annotations in proteins and design a multi-level alignment module to align the representations of sequences and annotations extracted from the existing encoders in both global and local level. For the generation task, PAAG utilizes an autoregressive decoder to generate protein sequences guided by the aligned representation of textual annotations. Additionally, PAAG employs an end-to-end training pipeline that joint training of alignment and generation tasks without freezing the parameters in sequence and text encoders. This joint training enhances the understanding of the complex and flexible annotation condition, resulting in improved guided generation. Figure 1 demonstrates an example of annotation-guided generation. In experiments, we first investigate the quality of protein representations from PAAG by seven predictive downstream tasks. PAAG surpasses state-of-the-art baseline with an average relative improvement of 1.5%. Subsequently, three protein generation tasks are conducted to assess the capabilities of PAAG. In the case of unconditional generation, PAAG produces sequences exhibiting the highest degree of novelty while maintaining the distribution of natural proteins. For the two conditional protein design tasks, PAAG demonstrates a nearly threefold increase in generation success rate (24.7% vs 4.7% in zinc finger, and 54.3% vs 22.0% in the immunoglobulin domain) [^1] This is the success rate with quality threshold $e=1$, i.e., $\text{SR}_{1}$. in comparison to the existing model. Our contributions are summarized as follows:
@@
  • We propose the first annotation-guided protein design paradigm, integrating local and global annotation information. Our proposed framework PAAG is the first approach that can generate proteins containing specific domains, guided by their corresponding annotations with high success rate.
@@
  • Comprehensive experiments on 7 predictive and 3 generative tasks showcase PAAG’s superiority compared to the existing methods. Notably, PAAG is not only capable of generating proteins that include a single annotation, but it also successfully generates proteins adhering to the flexible conditions of multiple annotation combinations.
2. Preliminaries
2.1. Protein and Its Textual Annotations
@@
The primary structure of a protein can be represented as an amino acid sequence $S=(s_{1},s_{2},\cdots,s_{n})$, where $s_{i}$ is the $i$-th amino acids chosen from 20 different amino acids which are represented as 20 characters. Given a protein $S$, the annotation-sequence pair set $\mathcal{D}_{S}=\{(T_{k},S^{k}_{i:j})\}_{k=1}^{K_{S}}$ is constructed by extracting the correspondence between textual annotations and protein domains & properties from protein database, such as UniProtKB (Consortium, 2022), where $T_{k}$ is the textual annotations of the $k$-th domain $S^{k}_{i:j}$ of the protein and $K_{S}$ is the number of annotation-domain pairs of the protein $S$. For the property annotation, the corresponding domain is the entire sequence, i.e, $S_{1:|S|}$. This annotation-domain pair set $\mathcal{D}_{S}$ provides comprehensive knowledge about the protein $S$ and, therefore, plays an important role in training the protein generation model.
2.2. Encoders for Proteins and Annotations
@@
We adopt a protein encoder (PE) to generate both the protein and its domain representations based on the amino acid sequence as $\bm{z}_{S}=f_{\text{PE}}(S)$, where $f_{\text{PE}}$ is the protein encoder and $\bm{z}_{S}$ is the embedding of $S$. For textual annotations, we generate their representations using the pre-trained language model $f_{\text{LM}}$ as $\bm{z}_{\mathcal{A}_{S}}=f_{\text{LM}}(G(\mathcal{A}_{S}))$, where $\mathcal{A}_{S}\subseteq\{T_{k}|(T_{k},S_{i:j}^{k})\in\mathcal{D}_{S}\}$ is a subset of the annotations in the annotation-domain pair set for a protein $S$, $G()$ is a template function which converts a set of annotations to a textual description, the details of $G()$ can be found in Appendix A.2, and $\bm{z}_{\mathcal{A}}$ is the embedding of the annotation set. Note that the annotation subset can cover all annotations in $\mathcal{D}_{S}$ or just one. If $\mathcal{A}_{S}$ covers only one annotation $T_{0}$, we can skip the template function and directly generate the embedding of $T_{0}$ using $f_{\text{LM}}$.
@@
We employ transformer-based protein encoder and text encoder as our base model and initialize them with pre-trained models. Specifically, we employ SciBERT(Beltagy et al., 2019), which is pre-trained on computer science and biological datasets, to initialize the text encoder $f_{\text{LM}}$. The pre-trained protein encoder, ProtBERT(Elnaggar et al., 2021), is used to initialize the protein encoder $f_{\text{PE}}$.
@@
Figure 2. The overall framework of PAAG. The same parameters share the same color. PAAG contains three modules. (1) Protein & Annotation Encoding module encode the input protein sequence & domains and corresponding annotations to the embeddings. (2) Multi-level alignment module projects the protein and annotation embeddings into and employs Annotation-Protein Contrasive (APC) loss, Annotation-Domain Contrasive (ADC) loss and Annotation-Protein Matching (APM) loss to align them in a same latent space. (3) Conditional Protein Decoding accepts the annotation embedding as input and generate the protein sequence.
3. Methodology
@@
In this section, we propose a novel framework, Protein-Annotation Alignment Generation(PAAG), which enables the flexible annotations-guided protein design. Figure 2 illustrates the overall framework of PAAG. In the following, we present the details of each component of PAAG.
3.1. Multi-level Protein and Annotation Alignment
@@
We will first employ template function $G$ to translate these annotations as textual descriptions and then utilize a language model to extract the representation of the annotation set $\bm{z}_{\mathcal{A}}=f_{\text{LM}}(G(\mathcal{A}))$. Then the protein sequence is generated using a decoder $f_{D}$ based on this representation as $S=f_{\text{D}}(\bm{z}_{\mathcal{A}})$.
@@
(1) $\vspace{-1ex}s(\mathcal{A},S)=\frac{\langle\text{Proj}_{\mathcal{A}}(z_{\mathcal{A}}),\text{Proj}_{S}(\bm{z}_{S})\rangle}{\|\text{Proj}_{\mathcal{A}}(\bm{z}_{\mathcal{A}})\|\|\text{Proj}_{S}(\bm{z}_{S})\|},$
@@ 3.1.1. Local Alignment
Annotation-Domain Contrastive(ADC) Loss: In designing this loss, since our objective is on identifying functional regions within proteins, we consider amino acid sequences from the same protein as negative samples. Specifically, we adopt InfoNCE loss as the ADC loss to align the representations of local annotation $\bm{z}_{T}$ and functional domain $\bm{z}_{S_{i:j}}$ as
(2) $\vspace{-1ex}\begin{split}\mathcal{L}_{\mathrm{ADC}}=-\frac{1}{K_{S}}\sum_{k=1}^{K_{S}}&\log\frac{\exp(s({T_{k}},{S^{k}_{i:j}})/\tau)}{\sum_{n=1}^{N}\exp(s({T_{n}},{S^{n-}_{i:j}})/\tau)}\end{split},$
@@
3.1.2. Global Alignment
To enable global alignment, we utilize template function $G(\cdot)$ to construct protein-level textual description and form the positive protein-description pairs as $\{(G(\mathcal{A}_{S}),S)\}$. Since multiple annotations and the entire protein will be more complex, to enlarge the number of negative samples for global alignment, we follow the setting of MOCO (He et al., 2020) to construct momentum encoders $f_{m}$ as:
(3) $f_{m}\leftarrow mf_{m}+(1-m)f.$
where the encoder $f$ can be either protein encoder $f_{\text{PE}}$ or text encoder $f_{\text{LM}}$, and $m$ is the momentum hyperparameter. We follow implementation details in (Li et al., 2021) and (Li et al., 2022) to construct the momentum encoders for both encoders. The momentum encoders extract consistent features to increase the number of negative samples, and the dynamic dictionaries will store these features.
@@ 3.1.2. Global Alignment
(4) $\vspace{-1ex}p_{m}^{\mathrm{s2a}}(S)=\frac{\exp(s(A_{m},S)/\tau)}{\sum_{m=1}^{M}\exp(s(A_{m},S)/\tau)},$
(5) $\vspace{-0.5ex}p_{m}^{\mathrm{a2s}}(A)=\frac{\exp(s(A,S_{m})/\tau)}{\sum_{m=1}^{M}\exp(s(A,S_{m})/\tau)},$
@@
(6) $\vspace{-1ex}\begin{split}\mathcal{L}_{\mathrm{APC}}=\frac{1}{2}\mathbb{E}_{(S,A)\sim D}\big{[}\mathrm{H}(\vec{y}^{\mathrm{~{}s2a}}(S),\vec{p}^{\mathrm{~{}s2a}}(S))\\ +\mathrm{H}(\vec{y}^{\mathrm{~{}a2s}}(A),\vec{p}^{\mathrm{~{}a2s}}(A))\big{]}\end{split}.$
Through APC loss, PAAG aligns the representations of two modalities, improving the quality of aligned representation, which is crucial for downstream tasks such as classification and regression.
Annotation-Protein Matching (APM) Loss. Inspired by  (Li et al., 2022), we introduce a multi-modal encoder $f_{\text{ME}}$, which integrates the representation of annotations and protein as $\bm{z}_{S,\mathcal{A}}^{M}=f_{\text{ME}}(S,G(\mathcal{A}))$, to identify whether the given pairs of protein description $G(\mathcal{A})$ and protein $S$ are matched or not. Specially, the multimodal encoder $f_{\text{ME}}$ is transformer-based and shares the parameters of self-attention and feed-forward layers with $f_{\text{PE}}$ and has additional cross-attention layers between self-attention layers and feed-forward layers to integrate the text information. The cross-attention layers share the same cross-attention parameters as the decoder.
@@
(7) $\vspace{-1ex}\mathcal{L}_{\mathrm{APM}}=\mathbb{E}_{(A,S)}\mathrm{H}(\vec{y}^{\textrm{APM}}(A,S),\vec{p}^{\textrm{APM}}(A,S)).$
APM loss trains the cross-attention layers for integrating two modalities. These cross-attention will be shared with protein decoder, which helps protein decoder interpret textual information. In Sec. 4.7, we demonstrate the importance of APM loss.
3.2. Conditional Protein Decoding
Protein Decoder. Our ultimate goal is generating functional proteins conditioned on a set of annotations $\mathcal{A}=\{T_{k}\}_{k=1}^{K}$. We adopt an auto-regressive protein decoder $f_{\text{D}}$ that can receive the condition from the language model. Specifically, the decoder will first process the protein sequence through causal attention layers to enable auto-regressive generation, then integrate the information from annotations via cross-attention layers followed by feed-forward layers. Note that the causal attention layers are initialized using the same weights as the self-attention layers in the protein encoder and the feed-forward layers share the same parameters as the protein encoder. Here, the parameter-sharing mechanism enables higher training efficiency (Li et al., 2023). The cross-attention layers are randomly initialized and trained from scratch.
(8) $\vspace{-1.5ex}\mathcal{L}_{PM}=-\sum_{i=1}^{l}\log{p(S_{i}|S_{<i};\mathcal{A})},$
3.3. Training Objectives
In the training of PAAG, we optimize four objective functions. These functions are designed to align the representations between annotations and protein sequences, integrate the two modalities, and reconstruct the protein sequence. In contrast to ProteinDT (Liu et al., 2023a), which splits the training process into three separate stages, PAAG jointly optimizes these four objective functions in an end-to-end manner. The overall pretraining objective of PAAG is :
(9) $\vspace{-1ex}\mathcal{L}=\mathcal{L}_{\mathrm{ADC}}+\mathcal{L}_{\mathrm{APC}}+\mathcal{L}_{\mathrm{APM}}+\mathcal{L}_{\mathrm{PM}}.$
3.4. Annotation-guided Protein Design